What is the Bond Order of Li(2)(minus) anion molecule?
i am getting as .5 but the answer is 1.5.Please explain How?

Grade:11

5 Answers

pvn s
11 years ago

Li- has three electrons in valence shell (bonding electrons =3)

therfore, bond order =(nof of bonding electrons)/2. =3/2=1.5 (NO OF ANTI BONDING electrons =0)

Upvote this if it helps :D

pvn s
11 years ago

no of bonding elecrtons =3 (valence), anti bonding =0 (given its state)

therfore =bond order =3/2=1.2

mohit yadav
11 years ago

 the ans. is o.5 only 1.5 is the wrong ans.

shashank singhal
11 years ago

my answer is alo .5.

Akash Kumar Dutta
11 years ago

to those whom i have disapproved,,,its because valence electrons of Li=1...it is not 3!